A possible means for making an airplane invisible to radar is to coat the plane with an antireflective polymer. If radar waves have a wavelength of 3.40 cm and the index of refraction of the polymer is n = 1.80, how thick (in cm) would you make the coating? (Assume that the index of refraction of the plane is higher than that of the coating. Also assume that the radar waves are normal to the surface of the coating. Give the minimum thickness that would make the airplane invisible to radar.)

0.472 cm

What If? Newer radar systems now use the VHF and UHF bands in order to detect stealthy aircraft. If a radar system operates with a frequency of 519 MHz (in the UHF band), what minimum thickness of coating (in cm) is needed to render an aircraft invisible to this radar band?

  cm
Answer #1

Wavelength, λ = v/f = 3*10^8 m/s /519*10^6 Hz = 0.578 m = 57.8 cm

thickness for destructive interference, t = λ/4n = 57.8/4*1.8
t = 8.028 cm